# Relative entropy of squeezed states in Quantum Field Theory

### v1

@@ Abstract
Utilizing the Tomita-Takesaki modular theory, we derive a closed-form analytic expression for the Araki-Uhlmann relative entropy between a squeezed state and the vacuum state in a free relativistic massive scalar Quantum Field Theory within wedge regions of Minkowski spacetime. Similarly to the case of coherent states, this relative entropy is proportional to the smeared Pauli-Jordan distribution. Consequently, the Araki-Uhlmann entropy between a squeezed state and the vacuum satisfies all expected properties: it remains positive, increases with the size of the Minkowski region under consideration, and decreases as the mass parameter grows.
@@ I Introduction
Quantum information theory has emerged as a powerful framework for analyzing fundamental aspects of quantum mechanics [1], particularly within the context of relativistic quantum field theory (QFT) [2, 3, 4]. A central concept in this framework is the quantum relative entropy, which quantifies the distinguishability between quantum states and plays a crucial role in quantum statistical mechanics, thermodynamics (see, e.g., [5, 6]), and the study of entanglement in high-energy physics [7, 3].
@@
In this work, we focus on the Araki-Uhlmann relative entropy [8, 9, 10], a generalization of quantum relative entropy to the setting of von Neumann algebras. This quantity is particularly well-suited for relativistic QFT, where the algebraic approach provides a natural description of quantum states in different spacetime regions [11, 12]. Specifically, we study the relative entropy for a squeezed vacuum state, a class of states that play a key role in quantum optics [13], Bose-Einstein condensates [14] and QFT, particularly in phenomena such as particle production in curved spacetime [15, 16], the Unruh effect [17, 18, 19], relativistic Quantum Information [20], Quantum Chromodynamics [21] as well as in the study of the Bell-CHSH inequality [22, 23, 24, 25].
@@
A key property we shall rely on is that a squeezed state can be obtained from the vacuum state via a unitary transformation. This feature, when combined with the modular theory of Tomita-Takesaki [26, 3] and the Bisognano-Wichmann results [27, 28], allows us to evaluate the Araki-Uhlmann relative entropy explicitly. Indeed, the Tomita-Takesaki modular theory provides a rigorous mathematical framework for understanding modular automorphisms in QFT, while the Bisognano-Wichmann theorem characterizes the modular structure of wedge-localized quantum field theories in Minkowski spacetime. By leveraging these tools, we derive a closed-form analytical expression for the relative entropy of a squeezed vacuum state in a free relativistic massive scalar QFT and analyze its behavior in wedge regions of Minkowski spacetime.
@@
Our main result is that the Araki-Uhlmann relative entropy between a squeezed state and the vacuum state turns out to be proportional to the smeared Pauli-Jordan distribution. As a consequence, the relative entropy exhibits the same fundamental properties already found in the case of coherent states [3, 29, 30, 31, 32]: it remains positive, increases with the size of the Minkowski region under consideration, and decreases as the mass parameter grows. These findings provide further insight into the structure of entanglement and information-theoretic aspects of QFT, reinforcing the role of relative entropy as a key quantity in the study of quantum correlations in relativistic settings.
@@
The paper is structured as follows: Sec. II provides the general properties of squeezed states for a relativistic scalar QFT. In Sec. III, we review the necessary background on the modular theory of Tomita-Takesaki as applied to the Araki-Uhlmann relative entropy. In Sec. IV, we present the computation of the relative entropy between a squeezed state localized in the right wedge and the vacuum state. Sec. V collects our conclusions. Finally, Appendices A and B summarize key notions of the canonical quantization of the massive scalar field and provide additional technical details about squeezed states.
II Properties of the squeezed states in relativistic scalar Quantum Field Theory
@@
=${\cal M}=\{\;W_{g}=e^{i\varphi(g)}\,,\;supp(g)\subset{\cal W}_{R}\;\}^{"}\;,$ (1)
@@
As a result, the squeezed state $|\Psi\rangle$ takes the form
@@
Equations (19) and (20) indicate that the squeezing operator ${\cal S}_{f}$ induces an automorphism of the von Neumann algebra ${\cal M}$ [30]. As a result, the squeezed state $|\Psi\rangle$ is both cyclic and separating [30], a crucial property for computing the relative entropy between $|\Psi\rangle$ and the vacuum state $|\Omega\rangle$. This computation will be the focus of the next section.
@@ III The Araki-Uhlmann relative entropy and the Tomita-Takesaki modular theory
A particularly useful aspect of analyzing the Araki-Uhlmann relative entropy between a squeezed state and the vacuum state lies in the remarkable relation [30, 31]
@@
The validity of Eq. (27) stems from the unitarity of the squeezing operator ${\cal S}_{f}$. To see this explicitly, consider
@@ IV Evaluation of the Araki-Uhlmann relative entropy
$S(\Psi|\Omega)=-\frac{1}{2i}\frac{d}{ds}\left(\bra{\Psi}(\Delta_{\Psi|\Omega}^{is}-c.c.)\ket{\Psi}\right)\Big{|}_{s=0}=-\frac{\alpha_{f}^{2}{\cal C}_{f}^{2}}{(1-4\alpha_{f}^{2}||f||^{4})^{3/2}}||f||^{2}\Delta_{PJ}\left(f,f^{\prime}_{s}\Big{|}_{s=0}\right)\;.$ (40)
@@
Expression (40) represents the central result of this work, providing a closed-form analytic expression for the relative entropy between the states $(|\Psi\rangle,|\Omega\rangle)$.
At this stage, two observations can be made:
  • As in the case of coherent states [29, 30, 31, 32], the Araki-Uhlmann relative entropy between a squeezed state $|\Psi\rangle$ and the vacuum state $|\Omega\rangle$ is entirely determined by the smeared Pauli-Jordan distribution.
  • This property leads to the expected behavior of the relative entropy [29, 30, 31, 32]: i) it is positive, ii) it decreases as the mass parameter increases, and iii) it increases with the size of the space-time region considered.
@@ V Conclusion
In this work, we have derived a closed-form expression for the Araki-Uhlmann relative entropy between a squeezed state and the vacuum state in a free massive scalar QFT. Utilizing the framework of Tomita-Takesaki modular theory and the Bisognano-Wichmann theorem, we showed that the modular flow associated with the squeezed state coincides with that of the vacuum, a consequence of the unitarity of the squeezing operator. This key observation enabled us to compute the relative entropy analytically, revealing that it is governed entirely by the smeared Pauli-Jordan distribution.
Our results generalize previous findings on coherent states and confirm that the relative entropy between a squeezed state and the vacuum state preserves its expected structural features: it is strictly positive, increases with the size of the space-time region under consideration, and decreases as the mass parameter grows. These behaviors further reinforce the interpretation of relative entropy as a robust measure of quantum distinguishability and information content in relativistic quantum field-theoretic settings.
The techniques and results presented here contribute to the broader understanding of quantum information measures in QFT and provide a concrete case where the intricate algebraic structures of QFT yield exact, physically meaningful quantities. Future investigations may extend these methods to more general classes of states or interacting theories, shedding further light on the interplay between algebraic modular theory and quantum information in relativistic systems.
@@ A.3 Physical interpretation and causality
Appendix B Some details about the squeezed states in Quantum Field Theory
In this appendix, we aim to proof that the squeezing operator $\mathcal{S}_{f}$ defines an automorphism for the von Neumann algebra $\mathcal{M}$. We start with the squeezing operator [^2] We have included a factor of 1/2 for convenience, but it does not spoil our proof.
@@ Appendix B Some details about the squeezed states in Quantum Field Theory
with $h=g+\gamma f.$ This completes the proof that the squeezing operator $\mathcal{S}_{f}$ defines an automorphism for the von Neumann algebra $\mathcal{M}$.

### v2

# Relative entropy of single-mode squeezed states in Quantum Field Theory

@@ Abstract
Utilizing the Tomita-Takesaki modular theory, we derive a closed-form analytic expression for the Araki-Uhlmann relative entropy between a single-mode squeezed state and the vacuum state in a free relativistic massive scalar Quantum Field Theory within wedge regions of Minkowski spacetime. Similarly to the case of coherent states, this relative entropy is proportional to the smeared Pauli-Jordan distribution. Consequently, the Araki-Uhlmann entropy between a single-mode squeezed state and the vacuum satisfies all expected properties: it remains positive, increases with the size of the Minkowski region under consideration, and decreases as the mass parameter grows.
@@ I Introduction
Quantum information theory has emerged as a powerful framework for analyzing fundamental aspects of quantum mechanics [1], particularly within the context of relativistic Quantum Field Theory (QFT) [2, 3, 4]. A central concept in this framework is the quantum relative entropy, which quantifies the distinguishability between quantum states and plays a crucial role in quantum statistical mechanics, thermodynamics (see, e.g., [5, 6]), and the study of entanglement in high-energy physics [7, 3] (see [8, 9, 10] for early works and [11, 12, 13, 14, 15, 16, 17, 18, 19] for recent accounts).
@@
In this work, we focus on the Araki-Uhlmann relative entropy [20, 21, 22], a generalization of quantum relative entropy to the setting of von Neumann algebras. This quantity is particularly well-suited for relativistic QFT, where the algebraic approach provides a natural description of quantum states in different spacetime regions [23, 24]. Specifically, we study the relative entropy for single-mode squeezed vacuum states, a class of states that play a pivotal role in quantum optics [25], Bose-Einstein condensates [26] and QFT. Single-mode squeezed states are particularly significant in phenomena such as particle production in curved spacetime [27, 28], the Unruh effect [29, 30, 31], relativistic Quantum Information [32], Quantum Chromodynamics [33] as well as in the study of the Bell-CHSH inequality [34, 35, 36, 37].
@@
A key property of single-mode squeezed states is that they can be obtained from the vacuum state via a unitary transformation defined by a single-mode squeezing operator. This feature, when combined with the modular theory of Tomita-Takesaki [38, 3] and the Bisognano-Wichmann results [39, 40], allows us to evaluate the Araki-Uhlmann relative entropy explicitly. Indeed, the Tomita-Takesaki modular theory provides a rigorous mathematical framework for understanding modular automorphisms in QFT, while the Bisognano-Wichmann theorem characterizes the modular structure of wedge-localized quantum field theories in Minkowski spacetime. By leveraging these tools, we derive a closed-form analytical expression for the relative entropy of a single-mode squeezed vacuum state in a free relativistic massive scalar QFT and analyze its behavior in wedge regions of Minkowski spacetime (see [14, 41] for recent numerical investigations of squeezed states in QFT).
@@
Our main result is that the Araki-Uhlmann relative entropy between a single-mode squeezed state and the vacuum state is proportional to the smeared Pauli-Jordan distribution. Consequently, the relative entropy exhibits the same fundamental properties observed in the case of coherent states [3, 42, 43, 44, 45]: it remains positive, increases with the size of the Minkowski region under consideration, and decreases as the mass parameter grows. These findings provide deeper insight into the structure of entanglement and information-theoretic aspects of single-mode squeezed states in QFT, reinforcing the role of relative entropy as a key quantity in the study of quantum correlations in relativistic settings.
@@
The paper is structured as follows: Sec. II provides the general properties of single-mode squeezed states for a relativistic scalar QFT. In Sec. III, we review the necessary background on the modular theory of Tomita-Takesaki as applied to the Araki-Uhlmann relative entropy. In Sec. IV, we present the computation of the relative entropy between a single-mode squeezed state localized in the right wedge and the vacuum state. Sec. V collects our conclusions. Finally, Appendices A and B summarize key notions of the canonical quantization of the massive scalar field and provide additional technical details about single-mode squeezed states.
II Properties of the single-mode squeezed states in relativistic scalar QFT
@@
=${\cal M}=\{\;W_{g}=e^{i\varphi(g)}\,,\;supp(g)\subset{\cal W}_{R}\;\}{"}\;,$ (1)
@@
As a result, the single-mode squeezed state $|\Psi\rangle$ takes the form
@@
Equations (20) and (21) indicate that the single-mode squeezing operator ${\cal S}_{f}$ induces an automorphism of the von Neumann algebra ${\cal M}$ [43]. As a result, the single-mode squeezed state $|\Psi\rangle$ is both cyclic and separating [43], a crucial property for computing the relative entropy between $|\Psi\rangle$ and the vacuum state $|\Omega\rangle$. This computation will be the focus of the next section.
@@ III The Araki-Uhlmann relative entropy and the Tomita-Takesaki modular theory
A particularly useful aspect of analyzing the Araki-Uhlmann relative entropy between a single-mode squeezed state and the vacuum state lies in the remarkable relation [43, 44]
@@
The validity of Eq. (28) stems from the unitarity of the single-mode squeezing operator ${\cal S}_{f}$. To see this explicitly, consider
@@ IV Evaluation of the Araki-Uhlmann relative entropy
$S(\Psi|\Omega)=-\frac{1}{2i}\frac{d}{ds}\left(\bra{\Psi}(\Delta_{\Psi|\Omega}^{is}-\textrm{c.c.})\ket{\Psi}\right)\Big{|}_{s=0}=-\frac{\alpha_{f}^{2}{\cal C}_{f}^{2}}{(1-4\alpha_{f}^{2}||f||^{4})^{3/2}}||f||^{2}\Delta_{PJ}\left(f,f^{\prime}_{s}\Big{|}_{s=0}\right)\;.$ (41)
@@
Expression (41) represents the central result of this work, providing a closed-form analytic expression for the relative entropy between the single-mode squeezed state $|\Psi\rangle$ and the vacuum state $|\Omega\rangle$.
At this stage, a few observations can be made: As in the case of coherent states [42, 43, 44, 45], the Araki-Uhlmann relative entropy between a single-mode squeezed state $|\Psi\rangle$ and the vacuum state $|\Omega\rangle$ is entirely determined by the smeared Pauli-Jordan distribution. Moreover, it is worth noting that expression (41) can be written as
$S(\Psi|\Omega)=\frac{2\,\alpha_{f}^{2}{\cal C}_{f}^{2}}{(1-4\alpha_{f}^{2}||f||^{4})^{3/2}}||f||^{2}S_{\rm coherent}\;,$ (44)
where $S_{\rm coherent}$ stands for the relative entropy between the coherent state $\text{e}^{i\varphi(f)}\ket{0}$ and the vacuum state, see Eq. (19) of [45]. As such, one can make use of the analysis done in [43] and express $S(\Psi|\Omega)$ as an integral of the smeared energy-momentum tensor, see Appendix A.3 of [43]. This property leads to the expected behavior of the relative entropy [42, 43, 44, 45]: i) it is positive, ii) it decreases as the mass parameter increases, and iii) it increases with the size of the space-time region considered.
@@ V Conclusion
In this work, we have derived a closed-form expression for the Araki-Uhlmann relative entropy between a single-mode squeezed state and the vacuum state in a free massive scalar QFT. Utilizing the framework of Tomita-Takesaki modular theory and the Bisognano-Wichmann theorem, we demonstrated that the modular flow associated with the single-mode squeezed state coincides with that of the vacuum, a consequence of the unitarity of the single-mode squeezing operator. This key observation enabled us to compute the relative entropy analytically, revealing that it is governed entirely by the smeared Pauli-Jordan distribution.
Our results extend previous findings on coherent states and confirm that the relative entropy between a single-mode squeezed state and the vacuum state preserves its expected structural features: it is strictly positive, increases with the size of the space-time region under consideration, and decreases as the mass parameter grows. These behaviors underscore the utility of relative entropy as a robust measure of quantum distinguishability and information content for single-mode squeezed states in relativistic quantum field-theoretic settings.
The techniques and results presented here enhance the understanding of quantum information measures in QFT, particularly for single-mode squeezed states, and provide a concrete example where the intricate algebraic structures of QFT yield exact, physically meaningful quantities. Future investigations may extend these methods to more general classes of states, such as multi-mode squeezed states or states in interacting theories, further elucidating the interplay between algebraic modular theory and quantum information in relativistic systems.
@@ A.3 Physical interpretation and causality
Appendix B Some details about the single-mode squeezed states in QFT
In this appendix, we aim to proof that the single-mode squeezing operator $\mathcal{S}_{f}$ defines an automorphism for the von Neumann algebra $\mathcal{M}$. We start with the single-mode squeezing operator [^3] We have included a factor of 1/2 for convenience, but it does not spoil our proof.
@@ Appendix B Some details about the single-mode squeezed states in QFT
with $h=g+\gamma f.$ This completes the proof that the single-mode squeezing operator $\mathcal{S}_{f}$ defines an automorphism for the von Neumann algebra $\mathcal{M}$.